Write an equation of the line that passes through the given point and has the givent slope. Use the equation to find two other points on the line. Check your solutions.
(-2,2); m = 3

Answer by checkley71(8403)

Y=3X+8 (TO FIND THE Y INTERCEPT X HAS TO BE =0 THUS WITH A SLOPE OF 3 WE NEED TO MULTIPLY 2*3=6 WHICH HAS TO BE ADDED TO THE Y COORDINATE OF 2 THUS WE HAVE 2+6=8 FOR THE Y INTERCEPT. IT LOOKS COMPLICATED BUT IF YOU USE GRAPH PAPER AND DRAW THE A LINE THROUGH POINTS(-2,2) WITH A SLOPE OF 4 YOU WILL SEE THAT IT WORKS)
IF Y=5 THEN X=-1 OR (-1,5)
IF X=-3 THEN Y=-3*3+8 OR Y=-9+8 OR Y=-1 (-3,-1)